OpenStudy (anonymous):

Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power

OpenStudy (mathteacher1729):

Can you write this using the equation editor? (click the little blue "Equation" button at the bottom left of the text box where you reply to posts.)

OpenStudy (anonymous):

@mathteacher1729

OpenStudy (anonymous):

\[5^{3/4} =\sqrt[4]{75}\]

OpenStudy (anonymous):

\[5^{2/3}=\sqrt[3]{25}\]

OpenStudy (anonymous):

do you understand ?

OpenStudy (mathteacher1729):

@markiaG - I can't access the link. That's probably because it's an image on a password-protected website. @jamaica25 I am fairly certain "5 to the 3 over 4 power" means \(\large 5^{3/4}\) and then " to the 2 over 3 power" means \(\large \left(5^{3/4}\right)^{2/3}\) but I'd like to see the problem exactly as written to remove any ambiguity.